OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 48π yd?

OpenStudy (anonymous):

A. d = 24 yd r = 12 yd B. d = 24 yd r = 48 yd C. d = 12 yd r = 24 yd D. d = 48 yd r = 24 yd

OpenStudy (aum):

What is the formula for the circumference of a circle?

OpenStudy (anonymous):

um dont know

OpenStudy (anonymous):

help will medal and fan

TheSmartOne (thesmartone):

Like @aum said \[C=2\pi \times r\]

TheSmartOne (thesmartone):

Now solve for r and remember \[D=2r\]

OpenStudy (anonymous):

??>???

OpenStudy (anonymous):

will medal and fan if will help me

OpenStudy (anonymous):

heloooooooooooooooo!!!!!!!!!!!!!!!!!!!!???????????? anone waaaaaaaaaa

TheSmartOne (thesmartone):

\[48\pi=2\pi \times r\] \[r=24\] \[d=2\times24\] \[d=48\]